Some bank robbers leave town, speeding at 71 mph. Twelve minutes later, the police give chase, traveling at 80 mph. How long will it take after the robbery for the police to overtake the robbers? (Round your answer to three decimal places.)
h=

Answer by mananth(16946)

robbers speed = 71 mph
police speed = 80 mph
catch up speed = 80-71 = 9 mph
12 minutes = 1/5 hour
in 1/5 hour the robbers travel 1/5 * 71 = 71/5 miles
..
time = (71/5 )/ 9
time = 71/5 * 1/9
time = 71/45
time = 1.578 hours